OpenStudy (anonymous):

Simplify 5^3√96 MULTIPLE CHOICE Choices: 20^3√6 10^3√12 7^3√12 2^3√12

geerky42 (geerky42):

\(\Large 5^3 \sqrt{96} \) or \(\Large 5 \sqrt[3]{96}\)?

OpenStudy (anonymous):

the 2nd one!

geerky42 (geerky42):

HINT: \(\large96 = 2^5 \cdot 3 \)

OpenStudy (anonymous):

i got that, but none of the choices have ^5

geerky42 (geerky42):

It's because you haven't simplify radical. Can you do it?

OpenStudy (anonymous):

10^3√12

geerky42 (geerky42):

Yes, perfect.

geerky42 (geerky42):

So is it clear to you now?

OpenStudy (anonymous):

yes! thanks lol

geerky42 (geerky42):

No problem, glad I helped.